# Entropy Rectifying Guidance for Diffusion and Flow Models

### v1

@@ B.2 I-ERG Impact on quality
B.3 Attention rectification mechanisms
In Table 6, we compare different attention rectification mechanisms form the litterature, including energy smoothing (from SEG) (Hong, 2024), identity mapping (from PAG) (Ahn et al., 2024) and temperature reduction at the denoiser level as is done in I-ERG. We use a guidance scale of $5.0$ and $\tau=0.2$. Our experiments indicate better performance when using low attention temperature when compared to both energy smoothing and identity mapping in the attention. The most notable difference being in density which is $12$ points higher for temperature reduction compared to energy smoothing and $3$ points higher than identity mapping. Similarly, temperature reduction achieves the best FID at $14.43$ which is $0.6$ points lower than energy smoothing and $2.2$ points better than identity mapping. Compared to energy smoothing, Identity mapping, as proposed in Ahn et al. (2024), achieves better precision, density, coverage and CLIPSCore than energy smoothing but still underperforms when compared to temperature reduction. Similar results can be observed qualitatively in Figure 17, temperature reduction generates images that showcase better realism in terms of the details of the image such as wooden textures on the floor in the first row, and details on the musician’s face and the background int he second row. Similarly, other qualitative results provide evidence supporting this effect, for example in Figure 12 and Figure 18 we observe more detailed backgrounds when using I-ERG than SEG.
@@ B.3 Attention rectification mechanisms
FID ()($\downarrow$)( ↓ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ ) Density ()($\uparrow$)( ↑ ) Coverage ()($\uparrow$)( ↑ ) CLIPSCore ()($\uparrow$)( ↑ )
Energy Smoothing $\sigma=+\infty$ $15.06$ $66.39$ ${\bf 39.24}$ $102.82$ $69.45$ $26.60$
Identity mapping $16.62$ $69.40$ $35.37$ $111.83$ $70.76$ $26.42$
Temperature reduction ${\bf 14.43}$ ${\bf 69.95}$ $36.57$ ${\bf 114.50}$ ${\bf 71.44}$ ${\bf 26.71}$
Figure 7: Percentage of points in the attention matric that have a high maximum probability. We observe three different regimes depending on the depth of the block. For blocks [0-5], the maximal association probability is high with more than $25\%$ of tokens having a maximum probability superior to $0.95$, For blocks 0-25, the association probability is rarely above $95\%$. Finally, the proportion of high certainty associations grows back for deeper layers in the model [25-38], reaching almost $40\%$ for the last layers $35-38$.
B.4 Attention Blocks
@@
Our results (for $t=0.3$) are summarized in Figure 7. We find a similar pattern across the different sampling steps. We observe three different regimes through the different layers of the model. While the first and last stage have a reasonably high matching probability, the middle stage has much smaller maximum probabilities and thus combines information from multiple tokens.
With this in mind, we posit that the energy landscape manipulations should only happen in the middle regime, as it would not overly harm the representations of the model. To validate this intuition, we perform a grid search over the range of blocks in which the energy manipulation is performed, we find a similar trend with the emergence of three different regimes, we find the best performance when applying the energy guidance on the middle blocks.
Figure 8: Choice of Rectified Attention Layers. Impact of the attention layers where the energy landscape is modified for I-ERG. (x-axis: start block, y-axis: end block). Operating ERG on middle layers seems to favorably impact performance, while including later layers results in harmful effects.
Figure 9: Influence of I-ERG kickoff threshold $\kappa$. Experiment conducted on our T2I model using Euler sampler with $50$ steps. We observe different trends for different metrics, FID and Recall are improved as $\kappa$ goes to $1$, precision and density on the other hand are improved when $\kappa$ goes to 00, CLIPSCore, VQAScore and coverage are optimal for $\kappa$ in the middle of the range $\kappa\in[0.2,0.6]$.
B.5 I-ERG kickoff threshold
In Figure 9, we perform a sweep over the kickoff threshold $\kappa$ for the denoiser-level energy guidance, with the temperature rescaling parameter set as $\tau_{i}=0.01$.
@@ B.5 I-ERG kickoff threshold
As reported in Astolfi et al. (2024), there is a natural trade-off between different facets of the generation i.e quality, consistency and diversity. While lower thresholds achieve the best precision and density (quality), higher thresholds notably improve coverage, recall and FID (diversity). On the other hand, CLIPScore and VQA (consistency) are maximized for a mid level threshold. This indicates that $\kappa$ can be tuned according to the task at hand to achieve suitable results, we find in our experiments that setting $\kappa\in[0.2,0.4]$ achieves the most balanced performance across different settings.
Task Text encoder $\kappa$ $l^{i}_{\text{min}}$ $l^{i}_{\text{max}}$ $\tau_{i}$ $l^{c}_{\text{min}}$ $l^{c}_{\text{max}}$ $\tau_{c}$ $\alpha$ $\gamma$ $K$
@@
T5-xxl 1 22 0.1
Appendix C Implementation
@@
I-ERG.  Changing standard attention mechanism with the energy-based variant defined in Algorithm 1 and applying this for attention blocks between $l_{\text{min}}$ and $l_{\text{max}}$ in the negative/unconditional prediction. Only applied if $t>\kappa$.
@@
In Table 7, we provide different hyperparameter sets used for our ERG method under different settings. We consider these values as the defaults used for our experiments unless specified otherwise.
@@
Appendix D Combining different methods
@@
FID Precision Recall Density Coverage CLIPSCore
CFG $35.85$ $40.09$ ${\bf 43.17}$ $41.11$ $10.95$ $24.71$
ERG ${\bf 25.82}$ ${\bf 68.59}$ $38.66$ ${\bf 114.67}$ ${\bf 24.29}$ ${\bf 27.71}$
Table 8: Effect of ERG on Stable Diffusion 3 model. We compare standard classifier-free guidance with our method on COCO-5K benchmark. We observe significant improvements across the different metrics.
Appendix E Reimplementation
We implement our method on the open-source stable diffusion 3 model, we use the sd3-medium from diffusers library. We refer to Table 7 for details about the choice of hyperparameters, additional qualitative examples on Figure 19 show significant improvements in terms of image quality. In Table 8, we report quantitative results on this model on COCO-5K benchmark. For this experiment we use the standard setup with a guidance scale of $5.0$ and Euler sampler with $28$ steps. We observe significant improvements for all metrics other than recall when using ERG compared to standard classifier-free guidance. More precisely, we observe significant boosts in FID $-10.03$ pts, density $+73.56$, coverage $+13.34$ and CLIPSCore $+3.0$.
Appendix F Additional qualitative results
@@
In Figure 10, we provide a comparison between calssifier-free guidance and Entropy Rectifying Guidance on three different prompts with varying levels of detail, we showcase the conditional diversity of the methods by providing samples with $5$ different random seeds for each prompt. For longer prompts (top row), we observe similar consistency but higher diversity in terms of background colors, the dog’s physical traits etc. Similarly, second row shows increased diversity with LPL in terms of colors, textures and the drawing that the cat is holding. For very short prompts (third row), we observe certain redundancies in the standard CFG generations while LPL is able to generate high quality samples showcasing higher diversity (imagining the horse in different situations, different angles for the picture, different time of day etc).
@@
In Figure 11, we provide a comparison between classifier-free guidance and I-ERG showcasing the improvements in image quality when generating samples with the same random seed. By setting $\tau_{i}=0.2$, the image semantics are not steered excessively compared to the CFG generations, resulting in comparable generations. We observe noticeable improvements in terms of high level details in the images when using I-ERG.
@@
In Figure 12, we provide comparisons between vanilla generation (no guidance), our implementations of SAG, PAG, SEG, and I-ERG for unconditional image generation. While PAG, and SEG are also able to improve image quality to satisfying levels, I-ERG provides an additional level of detail that is not present in the other methods (table texture on second row, background in the third and fifth row, water texture in fourth row and grass texture in the fifth row.
@@
In Figure 13, we provide qualitative comparisons between different methods for text-to-image generation. We compare our method with CFG, CADS and APG. We also provide samples generated when combining our method with either APG or CADS. Each column corresponds to a unique random seed that is used for the generations. We observe that both CFG and CADS tend to generate images that have a cartoonish style, APG on the other hand generates images that are more realistic looking. ERG alone significantly boosts image realism and results in less saturated images (pure white/black backgrounds for example). Mixing, ERG with either CADS or APG results in significant improvements in both image quality and diversity.
@@
In Figure 14, we provide a qualitative comparison between APG and APG + ERG, similar to Figure 1, we observe improvements in image quality and diversity. Also noticeable is a drift from unrealistic image styles (cartoonish) towards more realistic as can be seen in the astronaut and panda examples.
@@
In Figure 15, we showcase the effect of the image kickoff threshold $\kappa$ and denoiser attention temperature $\tau_{i}$ by interpolating both parameters in an unconditional generation experiment. We observe that lowering the attention temperature $\tau_{i}\rightarrow 0$ results in improved image quality. Similarly, lowering the image kickoff threshold $\kappa$ further improves image quality but steers the semantics of the image away from the vanilla generation ($\kappa=1.0,\tau_{i}=1$). However, as the image level threshold gets smaller $\kappa\rightarrow 0$, the semantics of the original image diverge from the vanilla generations, resulting in images with largely different semantics. We find $\kappa\in[0.2,0.4]$ and $\tau_{i}=0.01$ to work well in most cases.
@@
In Figure 16, we illustrate the effect of applying I-ERG on different layers of the denoiser. This experiment is conducted for class-conditional ImageNet-1K generations with our DiT-XL/2 model with 28 layers. We apply I-ERG to a different range of layers where three layers are involved each time. Applying I-ERG to either very early or late layers results in similar effects, with noticeable over-saturation and unnatural patterns that are overly simplistic as can be seen in the dog’s fur. On the other hand, applying I-ERG to middle layers results in improved image quality and sharper details. We find $l_{\text{min}}=12,l_{\text{max}}=15$ for the rest of the experiments to work well in this setup.
In Figure 18, we provide qualitative results comparing I-ERG with SEG and CFG for class-conditional generation on ImageNet at $512$ resolution, using a DiT-XL/2 model.
@@
Similarly to the unconditional and text-to-image cases, we find I-ERG to achieve higher image quality when compared CFG and our reimplementation of other guidance methods.
@@
Figure 14: Qualitative comparison showcasing the effect of Entropy Rectifying Guidance (ERG) on generation quality when combined APG.
@@
Figure 15: Effect of I-ERG main parameters. Interpolating the image kickoff threshold $\kappa$ on the vertical axis and denoiser attention temperature $\tau_{i}$ on the horizontal axis, we can see the effect on image quality. Unconditional generation experiment, $w_{\text{CFG}}=3.0$.
none $3-6$ $6-9$ $9-12$
@@
Figure 18: Qualitative comparison on ImageNet-1k@512. Samples from the same colum are generated using the same random seed, we compare ERG with CFG and SEG. Our method produces images of better quality. Images generated using Euler sampler with 50 steps.
An elephant crossing a river. A child wearing an apron and sunglasses posing in front of a blackboard with heart drawings and ’ERG’ written in it.
@@
Figure 19: Qualitative example of I-ERG on SD3-medium model compared to standard CFG. We apply our I-ERG on sd3-medium model. We observe better image details using I-ERG (kid’s face and details on the blouse in first row. Panda’s fur, reflections on the sunglasses in the second row.

### v2

@@ D.2 I-ERG impact on quality
D.3 Attention rectification mechanisms
In Table˜6, we compare different attention rectification mechanisms form the litterature, including energy smoothing (from SEG) (Hong, 2024), identity mapping (from PAG) (Ahn et al., 2024) and temperature reduction at the denoiser level as is done in I-ERG. We use a guidance scale of $5.0$ and $\tau=0.2$. Our experiments indicate better performance when using low attention temperature when compared to both energy smoothing and identity mapping in the attention. The most notable difference being in density which is $12$ points higher for temperature reduction compared to energy smoothing and $3$ points higher than identity mapping. Similarly, temperature reduction achieves the best FID at $14.43$ which is $0.6$ points lower than energy smoothing and $2.2$ points better than identity mapping. Compared to energy smoothing, Identity mapping, as proposed by Ahn et al. (2024), achieves better precision, density, coverage and CLIPScore than energy smoothing but still underperforms when compared to temperature reduction. Similar results can be observed qualitatively in Figure˜17, temperature reduction generates images that showcase better realism in terms of the details of the image such as wooden textures on the floor in the first row, and details on the musician’s face and the background in the second row. Similarly, other qualitative results provide evidence supporting this effect, for example in Figure˜12 and Figure˜18 we observe more detailed backgrounds when using I-ERG than SEG.
@@ D.3 Attention rectification mechanisms
FID $(\downarrow)$ Density $(\uparrow)$ Coverage $(\uparrow)$ CLIPSCore $(\uparrow)$
Energy smoothing $\sigma=+\infty$ $15.06$ $102.82$ $69.45$ $26.60$
Identity mapping $16.62$ $111.83$ $70.76$ $26.42$
Temperature reduction $\bf 14.43$} $\bf 114.50$} $\bf 71.44$} $\bf 26.71$}
Figure 7: Percentage of points in the attention matric that have a high maximum probability. We observe three different regimes depending on the depth of the block. For blocks [0-5], the maximal association probability is high with more than $25\%$ of tokens having a maximum probability superior to $0.95$, For blocks 0-25, the association probability is rarely above $95\%$. Finally, the proportion of high certainty associations grows back for deeper layers in the model [25-38], reaching almost $40\%$ for the last layers 353835-38.
D.4 Attention blocks
@@
Our results (for $t=0.3$) are summarized in Figure˜7. We find a similar pattern across the different sampling steps. We observe three different regimes through the different layers of the model. While the first and last stage have a reasonably high matching probability, the middle stage has much smaller maximum probabilities and thus combines information from multiple tokens.
With this in mind, we posit that the energy landscape manipulations should only happen in the middle regime, as it would not overly harm the representations of the model. To validate this intuition, we perform a grid search over the range of blocks in which the energy manipulation is performed in Figure˜8, we find a similar trend with the emergence of three different regimes, we obtain the best performance when applying the energy guidance on the middle blocks.
Figure 8: Choice of rectified attention layers. Impact of the attention layers where the energy landscape is modified for I-ERG. (x-axis: start block, y-axis: end block). Operating ERG on middle layers seems to performa favorably, while including later layers results in harmful effects.
Figure 9: Influence of I-ERG kickoff threshold $\kappa$. Experiment conducted on our T2I model using Euler sampler with $50$ steps. We observe different trends for different metrics, FID and Recall are improved as $\kappa$ goes to $1$, precision and density on the other hand are improved when $\kappa$ goes to 0, CLIPScore, VQAScore and coverage are optimal for $\tau$ in the middle of the range $\kappa\in[0.2,0.6]$.
D.5 I-ERG kickoff threshold
In Figure˜9, we perform a sweep over the kickoff threshold $\kappa$ for the denoiser-level energy guidance, with the temperature rescaling parameter set as $\tau_{i}=0.01$.
@@ D.5 I-ERG kickoff threshold
As also reported by Astolfi et al. (2024), there seems to be a natural trade-off between different facets of the generation i.e quality, consistency and diversity. While lower thresholds achieve the best precision and density (quality), higher thresholds notably improve coverage, recall and FID (diversity). On the other hand, CLIPScore and VQA (consistency) are maximized for a mid level threshold. This indicates that $\kappa$ can be tuned according to the task at hand to achieve suitable results, we find in our experiments that setting $\kappa\in[0.2,0.4]$ achieves the most balanced performance across different settings.
Task Text encoder $\tau$ $l^{i}_{\text{min}}$ $l^{i}_{\text{max}}$ $\tau_{i}$ $l^{c}_{\text{min}}$ $l^{c}_{\text{max}}$ $\tau_{c}$ $\alpha$ $\gamma$ $K$
@@
Flan-T5-XL 1 22 0.1
Text-to-image (SD3-medium) OpenCLIP-ViT/G 0.2 10 13 0.01 1 31 0.1 1.0 1.0 1.0
CLIP-ViT/L 1 12 0.1
Flan-T5-XL 1 22 0.1
Table 8: FID tuned metrics. Instead of considering Pareto optimal parameter set, we tune each of the baselines for optimal FID.
Method CFG ERG APG ERG+APG PAG SAG SEG CADS ERG+CADS AG
FID 5.25 4.93 6.62 5.24 7.00 5.28 6.72 5.12 5.00 7.11
D.6 Optimizing for FID
As shown by Astolfi et al. (2024), FID-optimal checkpoints often lie far from the Pareto front when jointly considering realism, consistency, and diversity. Dhariwal and Nichol (2021) further show that lower guidance scales minimize FID and recall, while higher guidance scales improve perceptual sharpness and structure, measured by IS and precision even though they increase FID.
Consequently, we tune different methods for FID alone, with results reported in Table˜8. CFG achieves its best FID at a guidance scale of 1.25, while ERG achieves its optimal FID of 4.93 using a guidance scale of 1.25 and an ERG scale of 2.0. This demonstrates that ERG is capable of outperforming prior guidance methods (e.g., SEG, APG) when optimized for FID alone, in addition to providing better trade-offs under broader evaluation criteria as can be seen in the Pareto fronts in Figure 2 of the main paper.
D.7 Additional Metrics
For better comparability, we extend the results in Table˜1 by adding Inception Score (IS), sFID, and Precision & Recall in Table˜9.
Table 9: Additional metrics. Comparison of guidance methods. Best results are in bold.
Method CFG APG CADS PAG SAG SEG AutoGuidance ERG ERG+APG ERG+CADS
Precision 65.71 66.34 66.42 66.42 64.97 61.22 61.48 70.92 69.50 72.72
Recall 43.95 45.15 45.75 43.94 47.98 36.88 34.60 41.43 50.25 38.89
sFID 26.53 20.31 18.20 27.21 28.55 33.80 14.75 19.56 9.30 15.32
IS 37.25 43.76 40.75 40.50 38.28 40.64 41.81 43.16 53.47 42.52
For precision & recall, we observe that similar trends to density and coverage with ERG+CADS achieving the best precision (72.72). ERG on its own achieves a precision of 70.92 compared to 65.71 for CFG. Recall is slightly lower when using ERG alone (41.43 vs. 43.95 for CFG), but ERG+APG achieves the best recall of 50.25. For Inception Score, ERG improves over the CFG baseline (43.16 vs. 37.25 for CFG) and achieves the best score when combined with APG (53.47). Simiwlarly, ERG+APG results in significant sFID and Inception score improvement, with ERG alone on par with APG.
Appendix E Implementation
@@
I-ERG.  Changing standard attention mechanism with the energy-based variant defined in Algorithm˜2 and applying this for attention blocks between $l_{\text{min}}$ and $l_{\text{max}}$ in the negative/unconditional prediction. Only applied if $t>\kappa$.
@@
In Table˜7, we provide different hyperparameter sets used for our ERG method under different settings. We consider these values as the defaults used for our experiments unless specified otherwise.
@@
Appendix F Combining different methods
@@
Appendix G Reimplementation
We implement our method on the open-source Stable Diffusion 3 model, we use the sd3-medium from diffusers library. We refer to Table˜7 for details about the choice of hyperparameters. Qualitative examples on Figure˜19 show significant improvements in terms of image quality. In Table˜10, we report quantitative results on this model on COCO-5k benchmark. For this experiment we use the standard setup with a guidance scale of $5.0$ and Euler sampler with $28$ steps. We observe significant improvements for all metrics when using ERG compared to standard classifier-free guidance. More precisely, we observe significant boosts in FID 10.03-10.03 pts, density +73.56+73.56, coverage +13.34+13.34 and CLIPScore +3.0+3.0.
Table 10: Effect of ERG on Stable Diffusion 3 models. We compare standard classifier-free guidance with our method on COCO-5K benchmark. We observe significant improvements across reported metrics.
FID $(\downarrow)$ Density $(\uparrow)$ Coverage $(\uparrow)$ CLIPScore $(\uparrow)$
SD3.0 CFG $35.85$ $41.11$ $10.95$ $24.71$
ERG $\bf 25.82$} $\bf 114.67$} $\bf 24.29$} $\bf 27.71$}
SD3.5 CFG $34.33$ $37.94$ $11.76$ $28.66$
ERG $\bf 23.81$} $\bf 118.59$} $\bf 25.26$} $\bf 29.61$}
PixART-alpha CFG $28.64$ $52.98$ $13.46$ $26.03$
ERG $\bf 27.75$} $\bf 70.61$} $\bf 16.57$} $\bf 26.61$}
SDXL CFG $21.65$ $80.93$ $20.00$ $29.54$
ERG $\bf 20.94$} $\bf 84.05$} $\bf 20.21$} $\bf 29.97$}
We implement ERG on different architectures beyond the standard DiT. Specifically, we provide additional (I-)ERG results for SDXL, PixArt-alpha and EDM2 (autoguidance). SDXL and EDM2 use U-Net architectures while PixArt-alpha uses local window attention. For SDXL and Pixart-alpha, we use the opensource version from Hugging Face alongside the EvalGIM evaluation framework. For EDM2 and DiT, we use the evaluation setups from their respective open-source implementations.
On EDM2/Autoguidance.  As reported in Table˜11, ERG improves over baseline EDM2 and obtains similar FID and slightly better FID_dinov2 than the autoguidance results, without need for additional models.
Table 11: EDM and Autoguidance comparison
Model/Metric FID $(\downarrow)$ FID_dinov2 $(\downarrow)$
EDM2-IN-XXL-512 2.00 59.34
EDM2-IN-XXL-512 + ERG 1.35 31.32
EDM2-IN-XXL-512 + Autoguidance 1.33 31.56
On open source DiT.  For completeness, we provide comparative results of our method applied to the official opensource implementation of DiT (Peebles and Xie, 2023). We re-ran the experiment using the guided-diffusion library to match exactly the evaluation setup from the DiT paper for 10 different random seeds. The results of the rectified experiment are reported in Table˜13 and Table˜13.
Table 12: Open source DiT results. Compared with the standard DiT.
Model/Metric FID sFID Precision Recall IS
DiT-XL/2-256 2.38 4.55 82.77 57.94 277.96
DiT-XL/2-256 + ERG 2.15 4.32 86.90 60.03 295.03
Table 13: Statistical analysis on FID. We report aggregate statistics over 10 different random seeds, minimum and maximum values are reported between brackets.
IS FID sFID Precision Recall
Baseline
[…]
[270.67270.67273.39273.39]
[…]
[2.312.312.482.48]
[…]
[4.414.414.554.55]
[…]
[82.4882.4882.7582.75]
[…]
[57.3457.3458.5058.50]
I-ERG
[…]
[289.18289.18297.84297.84]
[…]
[2.092.092.182.18]
[…]
[4.264.264.334.33]
[…]
[86.8286.8287.0687.06]
[…]
[59.2159.2160.6460.64]
For the baseline, we obtain an average FID of 2.38 which is higher than the one reported in the paper but is plausible given the variance of the results. Compared to the baseline, I-ERG still achieves a better FID (2.15 vs. 2.38), and given the statistics, also yields smaller variance on all metrics other than recall, showing more stable performance while varying random seeds.
Appendix H Assets
In Table˜14 we provide the links to the datasets and models used in our work and their licensing.
Table 14: Reference for the different assets used in our work.
COCO’14 https://www.cocodataset.org
ImageNet https://www.image-net.org
CC12M https://github.com/google-research-datasets/conceptual-12m
YFCC100M https://www.multimediacommons.org
Llama3-8b https://huggingface.co/meta-llama/Meta-Llama-3-8B
Flan-T5-XL https://huggingface.co/google/flan-t5-xl
Stable Diffusion 3 https://huggingface.co/stabilityai/stable-diffusion-3-medium
EvalGIM https://github.com/facebookresearch/EvalGIM
Appendix I Additional qualitative results
@@
In Figure˜10, we provide a comparison between classifier-free guidance and Entropy Rectifying Guidance on three different prompts with varying levels of detail, we showcase the conditional diversity of the methods by providing samples with $5$ different random seeds for each prompt. For longer prompts (top row), we observe similar consistency but higher diversity in terms of background colors, the dog’s physical traits etc. Similarly, the second row shows increased diversity with ERG in terms of colors, textures and the drawing that the cat is holding. For very short prompts (third row), we observe certain redundancies in the standard CFG generations while ERG is able to generate high quality samples showcasing higher diversity (depicting the horse against different backgrounds, different points of view, different time of day etc.).
@@
In Figure˜11, we provide a comparison between classifier-free guidance and I-ERG showcasing the improvements in image quality when generating samples with the same random seed. By setting $\tau_{i}=0.2$, the image semantics are not steered excessively compared to the CFG generations, resulting in comparable generations. We observe noticeable improvements in terms of high level details in the images when using I-ERG.
@@
In Figure˜12, we provide comparisons between vanilla generation (no guidance), our implementations of SAG, PAG, SEG, and I-ERG for unconditional image generation. While PAG, and SEG are also able to improve image quality to satisfying levels, I-ERG provides an additional level of detail that is not present in the other methods. See for example the table texture on second column, background in the third and fifth column, water texture in fourth column and grass texture in the fifth column.
@@
In Figure˜13, we provide qualitative comparisons between different methods for text-to-image generation using a single prompt and multiple seeds. We compare our method with CFG, CADS and APG. We also provide samples generated when combining our method with either APG or CADS. In each column the same random seed is used for all generations. We observe that both CFG and CADS tend to generate images that have a cartoonish style, APG on the other hand generates images that are more realistic. ERG alone significantly boosts image realism and results in less saturated images (pure white/black backgrounds for example). Mixing, ERG with either CADS or APG results in significant improvements in both image quality and diversity.
@@
In Figure˜14, we provide a qualitative comparison between APG and APG + ERG, similar to Figure˜1, we observe improvements in image quality and diversity. Also noticeable is a drift from unrealistic image styles (cartoonish) towards more realistic as can be seen in the astronaut and panda examples.
@@
In Figure˜15, we showcase the effect of the image kickoff threshold $\kappa$ and denoiser attention temperature $\tau_{i}$ by interpolating both parameters in an unconditional generation experiment. We observe that lowering the attention temperature $\tau_{i}\rightarrow 0$ results in improved image quality. Similarly, lowering the image kickoff threshold $\kappa$ further improves image quality but steers the semantics of the image away from the vanilla generation ($\kappa=1.0,\tau_{i}=1$). However, as the image level threshold gets smaller $\kappa\rightarrow 0$, the semantics of the original image diverge from the vanilla generations, resulting in images with largely different semantics. We find $\kappa\in[0.2,0.4]$ and $\tau_{i}=0.01$ to work well in most cases.
@@
In Figure˜16, we illustrate the effect of applying I-ERG on different layers of the denoiser. This experiment is conducted for class-conditional ImageNet-1k generations with our DiT-XL/2 model with 28 layers. We apply I-ERG to a different range of layers where three layers are involved each time. Applying I-ERG to either very early or late layers results in similar effects, with noticeable over-saturation and unnatural patterns that are overly simplistic as can be seen in the dog’s fur. On the other hand, applying I-ERG to middle layers results in improved image quality and sharper details. We find $l_{\text{min}}=12,l_{\text{max}}=15$ to work well and use it for the rest of the experiments in this setup.
@@
In Figure˜18, we provide qualitative results comparing I-ERG with SEG and CFG for class-conditional generation on ImageNet at $512$ resolution, using a DiT-XL/2 model. Similarly to the unconditional and text-to-image cases, we find I-ERG to achieve higher image quality when compared CFG and our reimplementation of other guidance methods.
@@
Figure 14: Qualitative comparison showcasing the effect of Entropy Rectifying Guidance (ERG) on generation quality when using APG.
@@
Figure 15: Effect of I-ERG main parameters. Interpolating the image kickoff threshold $\kappa$ on the vertical axis and denoiser attention temperature $\tau_{i}$ on the horizontal axis showcases the effect on image quality and global semantics of the generated image.
none 363-6 696-9 9129-12
@@
Figure 18: Qualitative comparison on ImageNet-1k@512. Samples from the same column are generated using the same random seed, we compare ERG with CFG and SEG. Our method produces images of better quality. Images generated using Euler sampler with 50 steps.
An elephant crossing a river.
A child wearing an apron and sunglasses posing in front of a blackboard with heart drawings and ’ERG’ written in it.
@@
A group of supporters wearing football jerseys cheering for their team after a victory.
A maltese dog getting a haircut by a hairdresser.
@@
Figure 19: Qualitative comparison using Stable Diffusion 3. For each prompt we use the same seed to sample images from sd3-medium (top rows) and sd3.5-large (bottom rows) with standard CFG (top) and I-ERG (bottom). We observe better image details using I-ERG.